OpenStudy (anonymous):

8x+9y=-15 Find the value of y when x=6.

OpenStudy (anonymous):

I should know these by now but I don't ;-;

OpenStudy (anonymous):

y=-7

OpenStudy (anonymous):

it can sometimes be forgettable

OpenStudy (anonymous):

8*6 +9y = -15 48 + 9y = -15 9y = -15-48 9y = -63 y = ?